# The Absolute Anabelian Geometry of Virtual Curves Arising from Sections of Arithmetic Fundamental Groups of Configuration Spaces

### v1

@@ 4. Weights of Determinant Maps
Let $G$ a profinite group; $l$ be a prime number; $M$, $N$ finitely generated $\hat{\mathbb{Z}}$-modules with continuous $G$-actions such that $\text{\rm Rank}_{l}(M)=n$, $\text{\rm Rank}_{l}(N)=1$. Suppose that we are given a $\hat{\mathbb{Z}}$-bilinear $G$-equivariant map $f:M\times M\rightarrow N$ (i.e., $f(g(a),g(b))=g(f(a,b))$ for any $a,b\in M$, $g\in G$) that induces an isomorphism $M\otimes_{\mathbb{Z}_{l}}\mathbb{Q}_{l}\simeq\text{\rm Hom}(M\otimes_{\mathbb{Z}_{l}}\mathbb{Q}_{l},N\otimes_{\mathbb{Z}_{l}}\mathbb{Q}_{l})$ of $\mathbb{Q}_{l}$-linear spaces. Moreover, suppose that the image of $\text{\rm Det}_{N}^{l}:G\to(\mathbb{Z}_{l}^{\times})^{\text{\rm tf}}$ is open. Then
@@
Let $k$ a field of characteristic zero; $X$ a smooth curve of type $(g,r)$ over $k$; $\bar{k}$ an algebraic closure of $k$; $l$ a prime number such that $X$ is $l$-cyclotomically full [cf. Definition 4.10(ii), where “$X$” is taken to be the $X$ of the present discussion]; $\text{\rm cycl}^{l}_{\Pi_{X}}:\Pi_{X}\to(\mathbb{Z}_{l}^{\times})^{\text{\rm tf}}$ the $l$-cyclotomic character; $\Delta_{X}\coloneqq\Pi_{X\times_{k}\bar{k}}$ the geometric fundamental group. Then
@@
For (ii), by Proposition 4.2(ii) and Proposition 4.6, it suffices to calculate $w^{l}_{\Pi_{X}/T(A)}$ where $A$ is the Albanese variety of $X$. Let
@@
of Tate modules. By Proposition 4.5 and the definition of Tate module, there exists a $l$-perfect pairing
@@
By definition, the action of $\Pi_{X}$ on $T(\mathbb{G}_{m})\otimes_{\mathbb{Z}_{l}}\mathbb{Q}_{l}\cong\mathbb{Q}_{l}$ is given by cyclotomic character, so $\text{\rm Det}^{l}_{T(\mathbb{G}_{m})}=\text{\rm cycl}^{l}_{\Pi_{X}}$. By Proposition 4.11, we have $w^{l}_{\Pi_{X}/T(A_{\text{\rm abel}})}=\text{\rm Rank}_{l}(T(A_{\text{\rm abel}}))$. Since $\text{\rm Rank}_{l}(T(A_{\text{\rm abel}}))=2g$ [cf. Proposition 4.2(i); [18], Remark 1.2.2], we obtain
@@
By Proposition 4.13, we may assume without loss of generality, by replacing $k$ by a suitable finite extension field of $k$, that $A_{\text{\rm tor}}$ is isomorphic to a product of copies of $\mathbb{G}_{m}$. Since the action of $\Pi_{X}$ on $T(\mathbb{G}_{m})\otimes_{\mathbb{Z}_{l}}\mathbb{Q}_{l}\cong\mathbb{Q}_{l}$ is given by cyclotomic character, by Proposition 4.12 and Proposition 4.2(ii), we conclude that
@@
Corollary 4.15 follows by combining Lemma 4.3, Theorem 4.14, and the fact that the cyclotomic character is functorial. ∎
@@
By [4], Exposé XII, Corollaire 5.2, for any positive integer $n$, there exists a cyclic étale cover of $X$ that is totally ramified at two cusps and unramified at the other cusps. By the Hurwitz formula, one sees that the type of this étale cover is $(ng,nr-2n+2)$. Now $\gcd(ng+nr-2n+1,g+r-1)=\gcd(1-n,g+r-1)$, so it suffices to choose an $n$ such that $n-1$ and $g+r-1$ is coprime. ∎
@@
By Corollary 4.15(ii), $w^{l}_{\chi}(\Delta_{\mathcal{C}^{\prime}})=2g(\mathcal{C}^{\prime})+2r(\mathcal{C}^{\prime})-2$, where $(g(\mathcal{C}^{\prime}),r(\mathcal{C}^{\prime}))$ is the type of $\mathcal{C}^{\prime}$. Let $\bar{y}$ be a geometric point of $Y$. By Proposition 3.7, there exists a natural isomorphism $\Pi_{X_{\bar{y}}}\xrightarrow{\sim}\Delta_{\mathcal{C}}$. By Lemma 4.16, there exists a set of Galois étale covers $\{Z_{i}\}_{i\in I}$ of $X_{\bar{x}}$, where $Z_{i}$ is of type $(g_{i},r_{i})$, such that the greatest common divisor of the elements of the set $\{g_{i}+r_{i}-1\}_{i\in I}$ is $1$. For each $Z_{i}$, write $\Delta_{i}$ for the corresponding open subgroup of $\Delta_{\mathcal{C}}$. For each $\Delta_{i}$, choose an open subgroup $\Pi_{i}\subset\Pi_{\mathcal{C}}$ such that $\Pi_{i}\cap\Delta_{\mathcal{C}}=\Delta_{i}$.
By Corollary 3.4(ii) and Corollary 3.6, for each $\Pi_{i}$, there exists a pointed virtual curve $\mathcal{C}_{i}:(X_{i},Y_{i},f_{i},k_{i},\bar{k},t_{i},s_{i})$ such that $\Pi_{\mathcal{C}_{i}}\simeq\Pi_{i}$. For each $Y_{i}$, choose an arbitrary geometric point $\bar{y}_{i}$ over $\bar{y}$. By definition, the type of $\mathcal{C}_{i}$ is equal to the type of $(X_{i})_{\bar{y}_{i}}$, which is isomorphic to $Z_{i}$ by construction. Proposition 4.17 now follows formally.
@@
First, we consider (i). Observe that $k$ is $l$-cyclotomically full if and only if the image of $\Pi_{\mathcal{C}}$ by $\text{\rm Det}^{l}_{\Delta_{\mathcal{C}}}$ is nontrivial. Choose an arbitrary $l$ that satisfies this purely group-theoretic condition. Then observe that for any group homomorphism $\chi:\Pi_{\mathcal{C}}\to(\mathbb{Z}_{l}^{\times})^{\text{\rm tf}}$ and any étale covering $\mathcal{C}^{\prime}$ of $\mathcal{C}$, $w^{l}_{\chi}(\Delta_{\mathcal{C}^{\prime}})$ may be recovered from the exact sequence in the statement of Proposition 4.18 and the image of $\Pi_{\mathcal{C}^{\prime}}$ in $\Pi_{\mathcal{C}}$. Therefore, by Proposition 4.17, the $l$-cyclotomic character $\Pi_{\mathcal{C}}\to(\mathbb{Z}_{l}^{\times})^{\text{\rm tf}}$ may be reconstructed from this exact sequence. By Corollary 4.15(ii)(iv), Corollary 3.4(ii) and Corollary 3.6, one may group-theoretically recover $2g(\mathcal{C^{\prime\prime}})+2r(\mathcal{C^{\prime\prime}})$ and $2g(\mathcal{C^{\prime\prime}})+r(\mathcal{C^{\prime\prime}})$ from the subgroup $H\subset\Pi_{\mathcal{C}}$ and this exact sequence. Therefore, by considering in addition whether or not $\Delta_{\mathcal{C}}$ is free, one may group-theoretically recover the type of $\mathcal{C^{\prime\prime}}$ from the subgroup $H\subset\Pi_{\mathcal{C}}$ and this exact sequence.
Next, we observe that (ii) follows immediately from Proposition 4.13, by applying the proof of (i) in the case where we take “$H$” to be $H\cap H^{\prime}$.
@@
(i) Let $\mathcal{C}:(X,Y,f,k,\bar{k},t,s)$ be a pointed virtual curve of type $(g,r)$ such that $k$ is of characteristic zero. Suppose that there exists a prime number $l$ such that $k$ is $l$-cyclotomically full. Then define the set of pre-cuspidal inertia subgroups to be the set of closed subgroups $I\subset\Delta_{\mathcal{C}}$ that are isomorphic to $\hat{\mathbb{Z}}$ and satisfy the condition that for any open subgroup $J\subset\Delta_{\mathcal{C}}$ such that $J\neq IJ$, $r(J)<[IJ:J]r(IJ)$ [cf. Proposition 4.18]. Define the set of quasi-cuspidal inertia subgroups to be the set of pre-cuspidal inertia subgroups $I\subset\Delta_{\mathcal{C}}$ such that every open subgroup of $I$ is a pre-cuspidal inertia subgroup. Define the set of cuspidal inertia subgroups to be the set of maximal elements in the set of quasi-cuspidal inertia subgroups with respect to the inclusion relation. Denote by $\text{\rm Cusp}(\mathcal{C})$ the set of conjugacy classes of cuspidal inertia subgroups in $\Delta_{\mathcal{C}}$.
(ii) Let $k$ be a field of characteristic zero, $X$ a smooth hyperbolic curve over $k$. Then denote by $\text{\rm Cusp}(X)$ the set of conjugacy classes of cuspidal inertia subgroups of $\Pi_{X}$, i.e., the intersections of decomposition groups of cusps with $\Delta_{X}$, the geometric fundamental group of $X$ [cf. § 2, the discussion entitled “Fundamental Groups”].
@@
First, we observe that (i) follows immediately from (ii) and (iii). (Here, we note that one may replace $k$ by a finite extension of $k$ without affecting the validity of (i). Therefore, one may assume, without loss of generality, that there exists a $k$-rational point $y$ of $Y$.) Moreover, (ii) follows immediately from Corollary 3.2 and Proposition 4.18.
@@
By Corollary 3.8, one may identify $\Pi_{\mathcal{C}}$ with $\Pi_{X_{y}}$. For any open subgroup $\square$ of $\Delta_{\mathcal{C}}$, denote by $X_{\square}$ the corresponding curve over $\bar{k}$. Thus by Proposition 4.18, $r(\square)=r(X_{\square})$.
Let $I$ be a quasi-cuspidal inertia subgroup of $\Delta_{\mathcal{C}}$. Let $p$ be a prime. Denote by $I_{p}$ the unique $p$-Sylow subgroup of $I$. Let $H$ be an open normal subgroup of $\Delta_{\mathcal{C}}$. We claim that the image of $I_{p}$ in $\Delta_{\mathcal{C}}/H$ is contained in the image of some cuspidal inertia group of $X_{y}$. Without loss of generality, one may assume that the image of $I_{p}$ in $\Delta_{\mathcal{C}}/H$ is nontrivial. By Definition 4.19(i), no nontrival intermediate subcovering (i.e., a covering $Z\to Z^{\prime}$ that appears in some factorization $X_{H}\to Z\to Z^{\prime}\to X_{HI_{p}}$) of the covering $X_{H}\to X_{HI_{p}}$ is unramified. Thus by considering the subcovering corresponding to $H\subset H(I_{p})^{p}\subset HI_{p}=HI_{p}$, we conclude (since $I_{p}$ is cyclic pro-$p$) that some cuspidal inertia subgroup of $X_{HI_{p}}$ surjects onto $HI_{p}/H$. Since $H$ is arbitrary, by [8], Lemma 1.5, we thus conclude that there exists a unique cuspidal inertia subgroup $J_{p}$ of $X_{y}$ such that $I_{p}\subset J_{p}$.
@@ 5. Characterization of Geometric Fundamental Groups
By Corollary 4.15(i), $\Delta_{\mathcal{C}}$ is a normal closed topologically finitely generated subgroup. Thus it suffices to show that for any normal closed topologically finitely generated subgroup $H$ of $\Pi_{\mathcal{C}}$, the image of $H$ in $G_{k}$ is trivial. But this follows from the fact that $G_{k}$ is very elastic, i.e., elastic and not topologically finitely generated [cf. [15], Theorem 1.7(iii)]. ∎
In the case of a mixed characteristic local field $k$, we will use Proposition 5.2 to calculate the absolute degree of $k$ from the abstract topological group $G_{k}$.
@@
Let $k$ be a mixed characteristic local field of residue characteristic $p$; $l\neq p$ a prime number. Denote by $[k:\mathbb{Q}_{p}]$ the dimension of $k$ over $\mathbb{Q}_{p}$. Then
@@
Let $l$ be a prime number, $k$ an $l$-cyclotomically full field of characteristic zero; $\mathcal{C}:(X,Y,f,k,\bar{k},t,s)$ a pointed virtual curve such that $Y$ is smooth over $k$. Suppose that there exist morphisms $u:Z\to Y$, $v:X\to Z$ where $u$ is proper, smooth, and geometrically connected of relative dimension one, and $v$ is an open immersion with dense image, such that $f=u\circ v$ [cf. § 2, the discussion entitled “Schemes and Curves”], and the reduced closed subscheme $D$ determined by the complement of the image of $X$ in $Z$ is (nessesarily finite) étale over $Y$. Let $D^{\prime}$ be a reduced closed subscheme of $Z$ that contains $D$ such that $D^{\prime}$ is (nessesarily finite) étale over $Y$. Denote by $X^{\prime}$ the complement of $D^{\prime}$ in $Z$. Write $i:X^{\prime}\to X$ for the natural open immersion; $f^{\prime}\coloneqq f\circ i$; $r$ for the degree of $D$ over $Y$; $r^{\prime}$ for the degree of $D^{\prime}$ over $Y$. Then the following properties hold:
@@
(ii) The induced maps $\pi_{\mathcal{F}}:\Pi_{\mathcal{C}^{\prime}}\rightarrow\Pi_{\mathcal{C}}$, $\delta_{\mathcal{F}}:\Delta_{\mathcal{C}^{\prime}}\rightarrow\Delta_{\mathcal{C}}$ are surjections. Moreover, the kernel of $\delta_{\mathcal{F}}$ and $\pi_{\mathcal{F}}$ are equal and generated by the subgroups in a collection of $r^{\prime}-r$ conjugacy classes of cuspidal inertia subgroups [cf. Definition 4.19].
@@
By Corollary 3.2 and Theorem 4.20(ii), the validity of (ii) is independent of the choice of section $s$. By Theorem 4.20(ii), one may replace $k$ by a finite field extension of $k$. Therefore, without loss of generality, one may assume that $s$ arises from a $k$-rational point $y$ of $Y$. By Corollary 3.8 and Theorem 4.20(iii), one may replace various virtual fundamental groups (and cuspidal inertia groups of virtual fundamental groups) by the étale fundamental groups of the respective fibers over $y$ (and the cuspidal inertia groups of the respective fibers over $y$). Then (ii) follows from well-known general facts concerning étale fundamental groups of curves.
Finally, (iii) follows from (ii), Lemma 4.3, Proposition 4.2(i)(ii) (where we note the natural functorial relationship between Proposition 4.2(i) and Proposition 4.2(ii)), and the fact that $k$ is $l^{\prime}$-cyclotomically full for arbitrary prime $l^{\prime}$ [cf. Proposition 5.2]. ∎
@@
(i) We refer to any morphism of pointed virtual curves that is isomorphic to some $\mathcal{F}$ of the sort constructed in Proposition 5.3 as a decuspidalization of pointed virtual curves.
(ii) Suppose that in the notation of Proposition 5.3, $D^{\prime}\setminus D$ is of degree $n$ over $Y$ (i.e., $n\coloneqq r^{\prime}-r$, where $r,r^{\prime}$ are defined in Proposition 5.3). Then we say that $\mathcal{F}$ is a degree $n$ decuspidalization. If $n=1$, then we refer to the unique conjugacy class of cuspidal inertia groups that appears in Proposition 5.3(ii) as the conjugacy class (of cuspidal inertia groups) determined by $\mathcal{F}$.
@@
We maintain the notation of Proposition 5.3. Suppose that $\mathcal{F}$ is a degree $1$ decuspidalization. Let $I\subset\Pi_{\mathcal{C}}$ be one of the representatives of the conjugacy class determined by $\mathcal{F}$. Assume further that $s$ arises from a $k$-rational point $y\in Y$. Then:
@@
Indeed, (i) follows from Corollary 3.8, while (ii) follows from (i) and the proof of Proposition 5.3(ii). ∎

### v2

@@ 4. Weights of Determinant Maps
Let $G$ a profinite group; $l$ be a prime number; $M$, NN finitely generated $^\hat{\mathbb{Z}}$-modules with continuous $G$-actions such that $\text{\rm Rank}_{l}(M)=n$, $\text{\rm Rank}_{l}(N)=1$. Suppose that we are given a $^\hat{\mathbb{Z}}$-bilinear $G$-equivariant map $f:M\times M\rightarrow N$ (i.e., $f(g(a),g(b))=g(f(a,b))$ for any $a,b\in M$, $g\in G$) that induces an isomorphism $M\otimes_{\mathbb{Z}_{l}}\mathbb{Q}_{l}\simeq\text{\rm Hom}(M\otimes_{\mathbb{Z}_{l}}\mathbb{Q}_{l},N\otimes_{\mathbb{Z}_{l}}\mathbb{Q}_{l})$ of $\mathbb{Q}_{l}$-linear spaces. Moreover, suppose that the image of $\text{\rm Det}_{N}^{l}:G\to(\mathbb{Z}_{l}^{\times})^{\text{\rm tf}}$ is open. Then
@@
Let $k$ a field of characteristic zero; $X$ a smooth curve of type $(g,r)$ over $k$; $\bar{k}$ an algebraic closure of $k$; $l$ a prime number such that $X$ is $l$-cyclotomically full [cf. Definition˜4.10(ii), where “XX” is taken to be the $X$ of the present discussion]; $\text{\rm cycl}^{l}_{\Pi_{X}}:\Pi_{X}\to(\mathbb{Z}_{l}^{\times})^{\text{\rm tf}}$ the $l$-cyclotomic character; $\Delta_{X}\coloneqq\Pi_{X\times_{k}\bar{k}}$ the geometric fundamental group. Then
@@
For (ii), by Proposition˜4.2(ii) and Proposition˜4.6, it suffices to calculate $w^{l}_{\Pi_{X}/T(A)}$ where $A$ is the Albanese variety of $X$. Let
@@
of Tate modules. By Proposition˜4.5 and the definition of Tate module, there exists a $l$-perfect pairing
@@
By definition, the action of $\Pi_{X}$ on $T(\mathbb{G}_{m})\otimes_{\mathbb{Z}_{l}}\mathbb{Q}_{l}\cong\mathbb{Q}_{l}$ is given by cyclotomic character, so $\text{\rm Det}^{l}_{T(\mathbb{G}_{m})}=\text{\rm cycl}^{l}_{\Pi_{X}}$. By Proposition˜4.11, we have $w^{l}_{\Pi_{X}/T(A_{\text{\rm abel}})}=\text{\rm Rank}_{l}(T(A_{\text{\rm abel}}))$. Since $\text{\rm Rank}_{l}(T(A_{\text{\rm abel}}))=2g$ [cf. Proposition˜4.2(i); [17], Remark 1.2.2], we obtain
@@
By Proposition˜4.13, we may assume without loss of generality, by replacing $k$ by a suitable finite extension field of $k$, that $A_{\text{\rm tor}}$ is isomorphic to a product of copies of $\mathbb{G}_{m}$. Since the action of $\Pi_{X}$ on $T(\mathbb{G}_{m})\otimes_{\mathbb{Z}_{l}}\mathbb{Q}_{l}\cong\mathbb{Q}_{l}$ is given by cyclotomic character, by Proposition˜4.12 and Proposition˜4.2(ii), we conclude that
@@
Corollary˜4.15 follows by combining Lemma˜4.3, Theorem˜4.14, and the fact that the cyclotomic character is functorial. ∎
@@
By [4], Exposé XII, Corollaire 5.2, for any positive integer $n$, there exists a cyclic étale cover of $X$ that is totally ramified at two cusps and unramified at the other cusps. By the Hurwitz formula, one sees that the type of this étale cover is (ng,nr2n+2)(ng,nr-2n+2). Now $\gcd(ng+nr-2n+1,g+r-1)=\gcd(1-n,g+r-1)$, so it suffices to choose an $n$ such that n1n-1 and g+r1g+r-1 is coprime. ∎
@@
By Corollary˜4.15(ii), $w^{l}_{\chi}(\Delta_{\mathcal{C}^{\prime}})=2g(\mathcal{C}^{\prime})+2r(\mathcal{C}^{\prime})-2$, where $(g(\mathcal{C}^{\prime}),r(\mathcal{C}^{\prime}))$ is the type of $\mathcal{C}^{\prime}$. Let $\bar{y}$ be a geometric point of $Y$. By Proposition˜3.7, there exists a natural isomorphism $\Pi_{X_{\bar{y}}}\xrightarrow{\sim}\Delta_{\mathcal{C}}$. By Lemma˜4.16, there exists a set of Galois étale covers $\{Z_{i}\}_{i\in I}$ of $X_{\bar{x}}$, where $Z_{i}$ is of type $(g_{i},r_{i})$, such that the greatest common divisor of the elements of the set $\{g_{i}+r_{i}-1\}_{i\in I}$ is $1$. For each $Z_{i}$, write $\Delta_{i}$ for the corresponding open subgroup of $\Delta_{\mathcal{C}}$. For each $\Delta_{i}$, choose an open subgroup $\Pi_{i}\subset\Pi_{\mathcal{C}}$ such that $\Pi_{i}\cap\Delta_{\mathcal{C}}=\Delta_{i}$.
By Corollary˜3.4(ii) and Corollary˜3.6, for each $\Pi_{i}$, there exists a pointed virtual curve $\mathcal{C}_{i}:(X_{i},Y_{i},f_{i},k_{i},\bar{k},t_{i},s_{i})$ such that $\Pi_{\mathcal{C}_{i}}\simeq\Pi_{i}$. For each $Y_{i}$, choose an arbitrary geometric point $\bar{y}_{i}$ over $\bar{y}$. By definition, the type of $\mathcal{C}_{i}$ is equal to the type of $(X_{i})_{\bar{y}_{i}}$, which is isomorphic to $Z_{i}$ by construction. Proposition˜4.17 now follows formally.
@@
$1\rightarrow\Delta_{\mathcal{C}}\rightarrow\Pi_{\mathcal{C}}\rightarrow\text{\rm Gal}(\bar{k}/k)\rightarrow 1.$
@@
First, we consider (i). Observe that $k$ is $l$-cyclotomically full if and only if the image of $\Pi_{\mathcal{C}}$ by $\text{\rm Det}^{l}_{\Delta_{\mathcal{C}}}$ is nontrivial. Choose an arbitrary $l$ that satisfies this purely group-theoretic condition. Then observe that for any group homomorphism $\chi:\Pi_{\mathcal{C}}\to(\mathbb{Z}_{l}^{\times})^{\text{\rm tf}}$ and any étale covering $\mathcal{C}^{\prime}$ of $\mathcal{C}$, $w^{l}_{\chi}(\Delta_{\mathcal{C}^{\prime}})$ may be recovered from the exact sequence in the statement of Proposition˜4.18 and the image of $\Pi_{\mathcal{C}^{\prime}}$ in $\Pi_{\mathcal{C}}$. Therefore, by Proposition˜4.17, the $l$-cyclotomic character $\Pi_{\mathcal{C}}\to(\mathbb{Z}_{l}^{\times})^{\text{\rm tf}}$ may be reconstructed from this exact sequence. By Corollary˜4.15(ii)(iv), Corollary˜3.4(ii) and Corollary˜3.6, one may group-theoretically recover $2g(\mathcal{C^{\prime\prime}})+2r(\mathcal{C^{\prime\prime}})$ and $2g(\mathcal{C^{\prime\prime}})+r(\mathcal{C^{\prime\prime}})$ from the subgroup $H\subset\Pi_{\mathcal{C}}$ and this exact sequence. Therefore, by considering in addition whether or not $\Delta_{\mathcal{C}}$ is free, one may group-theoretically recover the type of $\mathcal{C^{\prime\prime}}$ from the subgroup $H\subset\Pi_{\mathcal{C}}$ and this exact sequence.
Next, we observe that (ii) follows immediately from Proposition˜4.13, by applying the proof of (i) in the case where we take “HH” to be $H\cap H^{\prime}$.
@@
(i) Let $\mathcal{C}:(X,Y,f,k,\bar{k},t,s)$ be a pointed virtual curve of type $(g,r)$ such that $k$ is of characteristic zero. Suppose that there exists a prime number $l$ such that $k$ is $l$-cyclotomically full. Then define the set of pre-cuspidal inertia subgroups to be the set of closed subgroups $I\subset\Delta_{\mathcal{C}}$ that are isomorphic to $\hat{\mathbb{Z}}$ and satisfy the condition that for any open subgroup $J\subset\Delta_{\mathcal{C}}$ such that $J\neq IJ$, $r(J)<[IJ:J]r(IJ)$ [cf. Proposition˜4.18]. Define the set of quasi-cuspidal inertia subgroups to be the set of pre-cuspidal inertia subgroups $I\subset\Delta_{\mathcal{C}}$ such that every open subgroup of $I$ is a pre-cuspidal inertia subgroup. Define the set of cuspidal inertia subgroups to be the set of maximal elements in the set of quasi-cuspidal inertia subgroups with respect to the inclusion relation. Denote by $\text{\rm Cusp}(\mathcal{C})$ the set of conjugacy classes of cuspidal inertia subgroups in $\Delta_{\mathcal{C}}$.
(ii) Let $k$ be a field of characteristic zero, $X$ a smooth hyperbolic curve over $k$. Then denote by $\text{\rm Cusp}(X)$ the set of conjugacy classes of cuspidal inertia subgroups of $\Pi_{X}$, i.e., the intersections of decomposition groups of cusps with $\Delta_{X}$, the geometric fundamental group of $X$ [cf. §˜2, the discussion entitled “Fundamental Groups”].
@@
First, we observe that (i) follows immediately from (ii) and (iii). (Here, we note that one may replace $k$ by a finite extension of $k$ without affecting the validity of (i). Therefore, one may assume, without loss of generality, that there exists a $k$-rational point $y$ of $Y$.) Moreover, (ii) follows immediately from Corollary˜3.2 and Proposition˜4.18.
@@
By Corollary˜3.8, one may identify $\Pi_{\mathcal{C}}$ with $\Pi_{X_{y}}$. For any open subgroup $\square$ of $\Delta_{\mathcal{C}}$, denote by $X_{\square}$ the corresponding curve over $\bar{k}$. Thus by Proposition˜4.18, $r(\square)=r(X_{\square})$.
Let $I$ be a quasi-cuspidal inertia subgroup of $\Delta_{\mathcal{C}}$. Let $p$ be a prime. Denote by $I_{p}$ the unique $p$-Sylow subgroup of $I$. Let $H$ be an open normal subgroup of $\Delta_{\mathcal{C}}$. We claim that the image of $I_{p}$ in $\Delta_{\mathcal{C}}/H$ is contained in the image of some cuspidal inertia group of $X_{y}$. Without loss of generality, one may assume that the image of $I_{p}$ in $\Delta_{\mathcal{C}}/H$ is nontrivial. By Definition˜4.19(i), no nontrival intermediate subcovering (i.e., a covering $Z\to Z^{\prime}$ that appears in some factorization $X_{H}\to Z\to Z^{\prime}\to X_{HI_{p}}$) of the covering $X_{H}\to X_{HI_{p}}$ is unramified. Thus by considering the subcovering corresponding to $H\subset H(I_{p})^{p}\subset HI_{p}=HI_{p}$, we conclude (since $I_{p}$ is cyclic pro-$p$) that some cuspidal inertia subgroup of $X_{HI_{p}}$ surjects onto $HI_{p}/H$. Since $H$ is arbitrary, by [8], Lemma 1.5, we thus conclude that there exists a unique cuspidal inertia subgroup $J_{p}$ of $X_{y}$ such that $I_{p}\subset J_{p}$.
@@ 5. Characterization of Geometric Fundamental Groups
By Corollary˜4.15(i), $\Delta_{\mathcal{C}}$ is a normal closed topologically finitely generated subgroup. Thus it suffices to show that for any normal closed topologically finitely generated subgroup $H$ of $\Pi_{\mathcal{C}}$, the image of $H$ in $G_{k}$ is trivial. But this follows from the fact that $G_{k}$ is very elastic, i.e., elastic and not topologically finitely generated [cf. [14], Theorem 1.7(iii)]. ∎
In the case of a mixed characteristic local field $k$, we will use Proposition˜5.2 to calculate the absolute degree of $k$ from the abstract topological group $G_{k}$.
@@
Let $k$ be a mixed characteristic local field of residue characteristic $p$; $l\neq p$ a prime number. Denote by $k:\mathbb{Q}_{p}]$ the dimension of $k$ over $\mathbb{Q}_{p}$. Then
@@
Let $l$ be a prime number, $k$ an $l$-cyclotomically full field of characteristic zero; $\mathcal{C}:(X,Y,f,k,\bar{k},t,s)$ a pointed virtual curve such that $Y$ is smooth over $k$. Suppose that there exist morphisms $u:Z\to Y$, $v:X\to Z$ where $u$ is proper, smooth, and geometrically connected of relative dimension one, and $v$ is an open immersion with dense image, such that $f=u\circ v$ [cf. §˜2, the discussion entitled “Schemes and Curves”], and the reduced closed subscheme $D$ determined by the complement of the image of $X$ in $Z$ is (nessesarily finite) étale over $Y$. Let $D^{\prime}$ be a reduced closed subscheme of $Z$ that contains $D$ such that $D^{\prime}$ is (nessesarily finite) étale over $Y$. Denote by $X^{\prime}$ the complement of $D^{\prime}$ in $Z$. Write $i:X^{\prime}\to X$ for the natural open immersion; $f^{\prime}\coloneqq f\circ i$; $r$ for the degree of $D$ over $Y$; $r^{\prime}$ for the degree of $D^{\prime}$ over $Y$. Then the following properties hold:
@@
(ii) The induced maps $\pi_{\mathcal{F}}:\Pi_{\mathcal{C}^{\prime}}\rightarrow\Pi_{\mathcal{C}}$, $\delta_{\mathcal{F}}:\Delta_{\mathcal{C}^{\prime}}\rightarrow\Delta_{\mathcal{C}}$ are surjections. Moreover, the kernel of $\delta_{\mathcal{F}}$ and $\pi_{\mathcal{F}}$ are equal and generated by the subgroups in a collection of $r^{\prime}-r$ conjugacy classes of cuspidal inertia subgroups [cf. Definition˜4.19].
@@
By Corollary˜3.2 and Theorem˜4.20(ii), the validity of (ii) is independent of the choice of section $s$. By Theorem˜4.20(ii), one may replace $k$ by a finite field extension of $k$. Therefore, without loss of generality, one may assume that $s$ arises from a $k$-rational point $y$ of $Y$. By Corollary˜3.8 and Theorem˜4.20(iii), one may replace various virtual fundamental groups (and cuspidal inertia groups of virtual fundamental groups) by the étale fundamental groups of the respective fibers over $y$ (and the cuspidal inertia groups of the respective fibers over $y$). Then (ii) follows from well-known general facts concerning étale fundamental groups of curves.
Finally, (iii) follows from (ii), Lemma˜4.3, Proposition˜4.2(i)(ii) (where we note the natural functorial relationship between Proposition˜4.2(i) and Proposition˜4.2(ii)), and the fact that $k$ is $l^{\prime}$-cyclotomically full for arbitrary prime $l^{\prime}$ [cf. Proposition˜5.2]. ∎
@@
(i) We refer to any morphism of pointed virtual curves that is isomorphic to some $\mathcal{F}$ of the sort constructed in Proposition˜5.3 as a decuspidalization of pointed virtual curves.
(ii) Suppose that in the notation of Proposition˜5.3, $D^{\prime}\setminus D$ is of degree $n$ over $Y$ (i.e., $n\coloneqq r^{\prime}-r$, where $r,r^{\prime}$ are defined in Proposition˜5.3). Then we say that $\mathcal{F}$ is a degree $n$ decuspidalization. If $n=1$, then we refer to the unique conjugacy class of cuspidal inertia groups that appears in Proposition˜5.3(ii) as the conjugacy class (of cuspidal inertia groups) determined by $\mathcal{F}$.
@@
We maintain the notation of Proposition˜5.3. Suppose that $\mathcal{F}$ is a degree $1$ decuspidalization. Let $I\subset\Pi_{\mathcal{C}}$ be one of the representatives of the conjugacy class determined by $\mathcal{F}$. Assume further that $s$ arises from a $k$-rational point $y\in Y$. Then:
@@
Indeed, (i) follows from Corollary˜3.8, while (ii) follows from (i) and the proof of Proposition˜5.3(ii). ∎